OpenStudy (anonymous):

Last one Help mehh! :) ! The manager of Frank’s Snack Shop buys hot dogs in packages of 36. He buys hot dog buns in packages of 20. He cannot buy part of a package. What is the least number of packages of each product he can buy to have an equal number of hot dogs and buns? a. 5 packages of hot dogs and 9 packages of buns c. 10 packages of hot dogs and 10 packages of buns b. 9 packages of hot dogs and 5 packages of buns d. 12 packages of hot dogs and 8 packages of buns

OpenStudy (anonymous):

Rather then try to solve this one just cross check with the given options Option a is the answer Since 5*36 = 9*20 =180 :)

OpenStudy (anonymous):

:D!! thnx